OpenStudy (ffuentes6596):

Find the product of (3x − 4)(2x2 + 2x − 1). 6x3 + 2x2 − 5x + 4 6x3 + 14x2 − 11x + 4 6x3 − 14x2 − 5x + 4 6x3 − 2x2 − 11x + 4

OpenStudy (wanell):

(3x - 4)(2x^2 + 2x - 1) = 3x(2x^2 + 2x - 1) - 4(2x^2 + 2x - 1) = (6x^3 + 6x^2 - 3x) - (8x^2 + 8x - 4) = 6x^3 + 6x^2 - 3x - 8x^2 - 8x + 4 = 6x^3 - 2x^2 - 11x + 4

OpenStudy (wanell):

hope that helps

OpenStudy (ffuentes6596):

Thank you